OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3 y – 8 = 3 (x + 8) y + 8 = 3 (x + 8) y – 8 = 3 (x – 8) y + 8 = 3 (x – 8)

OpenStudy (anonymous):

ok so if you have a point (x1,y1) and a slope m the corresponding line equation is given by y = y1 + m(x-x1) can you do it now ?

OpenStudy (anonymous):

So, would it look like this y= -8+3(x-8) ?

OpenStudy (anonymous):

So, it would be the 1st one . Right?

OpenStudy (anonymous):

no .. notice that you have y = -8 +3(x-8) in order to bring it to the form of your options you have to add 8 to both sides what do you get ?

OpenStudy (anonymous):

I don't understand how to write it out ?

OpenStudy (anonymous):

y = -8 +3(x-8) add 8 to both sides : y + 8 = -8 +3(x-8) + 8 y + 8 = -8 + 8 + 3(x-8) y + 8 = 3(x-8)

jimthompson5910 (jim_thompson5910):

Another way to do it: y - y1 = m(x-x1) ... Start with the general point-slope form equation y - (-8) = 3(x-8) ... Plug in m = 3, and (x1, y1) = (8, -8) y + 8 = 3(x-8)

OpenStudy (anonymous):

Ohhhhh, ok thanks . Now I understand how to write it. Thank you for helping me! (: